OpenStudy (anonymous):

Please help! 1) Find the point on the curve \[y=x \sqrt{x}\] where the tangent line is parallel to the line 6x-y=4 2) At what point on the curve y=-2x^4 is the tangent line is perpendicular to the line x-y+1=0.

OpenStudy (anonymous):

i got the the first the question!

OpenStudy (anonymous):

what does it mean by perpendicular?

OpenStudy (anonymous):

@pooja195

ganeshie8 (ganeshie8):

Heyy still working on this ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

right now, i got y'=-8x^3

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{\normalsize \text{the slope of the line 6x-y=4 is}\hspace{.33em}\\~\\ y=6x-4 \hspace{.33em}\\~\\ m=6 \hspace{.33em}\\~\\ \normalsize \text{ For our tangent line to be parallel it must have the same slope. }\hspace{.33em}\\~\\ \normalsize \text{ So when is the slope of the tangent line 6 }\hspace{.33em}\\~\\ \normalsize \text{ Set the derivate equal to 6 }\hspace{.33em}\\~\\ y=x\sqrt{x} \hspace{.33em}\\~\\ y=x^{\frac{3}{2}} \hspace{.33em}\\~\\ y'=\dfrac{3}{2}x^{\frac{1}{2}} \hspace{.33em}\\~\\ 6=\dfrac{3}{2}x^{\frac{1}{2}} \hspace{.33em}\\~\\ x^{\frac{1}{2}}=4 \hspace{.33em}\\~\\ x=16 \hspace{.33em}\\~\\ \normalsize \text{But to find a point, we need x and y value. }\hspace{.33em}\\~\\ \normalsize \text{ So plug 16 in for your x value into your original equation }\hspace{.33em}\\~\\ y=6x-4 \hspace{.33em}\\~\\ y=6\times 16-4 \hspace{.33em}\\~\\ y=92 \hspace{.33em}\\~\\ \LARGE (x,y)\implies (16,92) \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

i think you sub in the wrong equation @mathmath333

OpenStudy (anonymous):

need help on the second question please. @ganeshie8